### v1

# Eirini Schoinas1, Adyah Rastogi2, Anissa Carter3, Jacob Granley2, and Michael Beyeler2,4

Abstract
Human-in-the-loop optimization (HILO) is a promising approach for personalizing visual prostheses by iteratively refining stimulus parameters based on user feedback. Previous work demonstrated HILO’s efficacy in simulation, but its performance with human participants remains untested. Here we evaluate HILO using sighted participants viewing simulated prosthetic vision to assess its ability to optimize stimulation strategies under realistic conditions. Participants selected between phosphenes generated by competing encoders to iteratively refine a deep stimulus encoder (DSE). We tested HILO in three conditions: standard optimization, threshold misspecifications, and out-of-distribution parameter sampling. Participants consistently preferred HILO-generated stimuli over both a naïve encoder and the DSE alone, with log odds favoring HILO across all conditions. We also observed key differences between human and simulated decision-making, highlighting the importance of validating optimization strategies with human participants. These findings support HILO as a viable approach for adapting visual prostheses to individuals.
I Introduction
Visual prostheses are being developed to restore vision for individuals with incurable blindness by electrically stimulating functional cells along the visual pathway [1, 2]. Retinal [3, 4, 5] and cortical [6, 7, 8] prostheses have enabled tasks such as object localization and supported mobility. However, the quality of vision provided by these implants remains limited. Phosphenes—the artificial visual percepts evoked by stimulation—vary widely across individuals [9, 10, 6, 11] and often do not combine linearly [12, 13, 14], suggesting the presence of a nonlinear transfer function between electrical stimuli and perceptual outcomes.
To address these nonlinearities, computational models have been developed to predict perceptual responses to electrical stimulation. These forward models use user-specific parameters to capture how stimulus properties affect the brightness and shape of phosphenes, as well as their interaction across electrodes [15, 9, 16, 17, 18]. While forward models enable a mechanistic understanding of prosthetic vision, their utility in real-world applications hinges on the ability to invert them: that is, to determine the optimal stimulus parameters required to elicit a desired percept. Deep stimulus encoders (DSEs) have been proposed for this purpose (Fig. 1A), leveraging deep learning to approximate the inverse mapping from percepts to stimuli [19, 20, 21, 17, 22]. However, DSEs require precise knowledge of user-specific parameters, which is often infeasible due to limited data availability or the inherent variability across users [23].
To optimize stimulus strategies from limited human feedback, the human-in-the-loop optimization (HILO) framework was introduced by Tristan and Fauvel [24] (Fig. 1B). Their work applied HILO to a linear phosphene model, demonstrating the feasibility of estimating parameters using pairwise comparisons of two competing simulated percepts (duels). Granley et al. [23] extended this approach by incorporating DSEs and testing it under more realistic conditions, such as model misspecifications and noisy user feedback. Their results suggested that HILO could efficiently adapt stimulus encoders to individual users, even when users selected duel outcomes randomly in two out of three duels. However, this work relied entirely on simulated users whose perceptual decisions followed predefined rules based on model predictions. Human decision making, by contrast, is inherently more variable and may not conform to model assumptions, leaving open the question of whether HILO would perform similarly when real users provide feedback.
Figure 1: A) Deep stimulus encoder (DSE). A forward model ($f$) predicts the perceptual response to visual stimuli based on user-specific parameters ($\phi$), while an encoder ($f^{-1}$) learns to minimize the perceptual error between predicted and target percepts. B) Human-in-the-loop optimization (HILO). The parameters from the DSE are refined using user preferences, collected through 60 binary comparison trials per condition. New parameter pairs are adaptively selected to efficiently converge on the most preferred percept. The target percept changes each iteration. Adapted under CC-BY from https://doi.org/10.48550/arXiv.2306.13104. C) Example duels used to infer user preferences are presented to sighted participants on a computer monitor. Participants selected the preferred stimulus based on both shape and brightness.
In this study, we take an intermediate step towards real-world validation by incorporating sighted participants viewing simulated prosthetic vision (SPV) stimuli on a monitor (Fig. 1C). This approach allows us to systematically test whether HILO can effectively adapt to individual perceptual variability, maintain robustness to model misspecifications, and remain practical under real-world decision-making constraints. By replacing simulated patient decisions with human feedback, this work provides critical evidence for the utility of HILO in personalizing neuroprosthetic devices.
II Methods
II-A Participants
Seventeen sighted undergraduate students from the University of California, Santa Barbara, participated in the study (10 female, 7 male; ages 18–21; $M=19.6$, $SD=1.12$). Participants were recruited through university-wide announcements and provided informed consent prior to participation. All participants reported normal or corrected-to-normal vision and no history of neurological or visual impairments. Participants were briefed on the study’s purpose and tasks before beginning the experiment. The study was approved by the Institutional Review Board at the University of California, Santa Barbara.
II-B Task
Participants viewed simulated prosthetic vision on a monitor and completed a series of trials aimed at optimizing the mapping between electrical stimuli and perceived visual representations. Each trial, referred to as a “duel,” presented two phosphenes generated for a target image. Participants selected the phosphene they perceived to better match the target image, described textually (e.g., “number eight”) rather than visually, to encourage independent mental representations of the target. To prevent selection bias, the positions of the two stimuli (left or right) were randomized.
In practice, electrical stimulation can evoke phosphenes that are excessively bright or even painful. To account for the limited dynamic range of computer monitors, numerical brightness values were displayed alongside each stimulus. During the tutorial phase, participants were introduced to the brightness scale: 1 represented threshold brightness (too dim), 2 was twice the threshold (ideal), and higher values indicated increasing excessive brightness (Fig. 2).
The experiment consisted of three phases:
  • Tutorial phase: Participants completed practice trials to familiarize themselves with the task, including how to make selections and interpret brightness values.
  • Optimization phase: Participants completed 60 duels, during which the HILO framework iteratively refined the simulated patient parameters $\phi$ based on their preferences.
  • Evaluation phase: Participants completed 39 additional duels, comparing the optimized DSE against a naïve baseline encoder.
Figure 2: Example stimuli illustrating the range of phosphene brightness levels shown to participants. A value of 0 represents complete darkness, 2 is the ideal brightness for a retinal prosthesis user, 5 is overly bright, and 10 is extremely bright, with white filling most of the stimulus area.
II-C Simulated Prosthetic Vision
We employed the experimentally validated computational model from [23] to replicate the perceptual experience of an epiretinal implant. This model represents phosphenes as multivariate Gaussian blobs, whose size, shape, and brightness are determined by both electrode location and stimulus properties such as amplitude, frequency, and pulse duration [25, 26, 9]. Importantly, the appearance of these phosphenes varies across individuals due to differences in retinal anatomy and neural processing.
To capture this variability, the model includes a set of 13 user-specific parameters, $\phi$, which govern key aspects of phosphene appearance. Among these, $\rho$ (in microns) controls overall phosphene size, while $\lambda$ (ranging from 0 for circular phosphenes to near 1 for highly elongated ones) modulates elongation along the trajectory of underlying axon pathways. Other parameters influence brightness scaling, the spread of axonal streaking effects, the location of the optic disc (which influences axonal trajectories), the implant location and orientation relative to the fovea, and electrode sensitivity. For full details, see [23]. Each participant was randomly assigned a unique set of $\phi$ values, simulating the perceptual variability observed in real prosthesis users.
The model maps an electrical stimulus $\mathbf{s}\in\mathbb{R}^{n_{e}\times 3}$ to a visual percept. Each phosphene is modeled as a Gaussian blob with center $\mu_{e}$, covariance matrix $\Sigma_{e}$, and brightness $b_{e}$, determined by both $\phi$ and the applied stimulus [23]:
$b(x,y)=2\pi b_{e}\det{(\mathbf{\Sigma_{e}}})\,\,\mathcal{N}([x,y]^{\top}|\mu_{e},\mathbf{\Sigma_{e}}),$ (1)
where $\mathcal{N}$ represents a Gaussian distribution, $\det{(\Sigma_{e})}$ ensures proper normalization of the percept brightness, and $\Sigma_{e}$ encodes phosphene shape and orientation, incorporating effects of retinal fiber structure.
To form the final percept, phosphenes from all stimulated electrodes are summed across the visual field. This summation introduces nonlinear interactions between adjacent phosphenes, which can lead to perceptual distortions similar to those reported by real prosthesis users [15].
This model serves as the foundation for generating the SPV stimuli used in this study.
II-D Human-in-the-Loop Optimization (HILO)
Following [23], a neural network DSE was trained to invert the forward model, generating stimulus parameters that elicit a phosphene most closely matching a target image. The DSE was designed as a fully connected feedforward network with multiple residual blocks, each containing batch normalization and leaky ReLU activations to stabilize training and improve generalization. The model takes as input the target image and the user-specific parameters $\phi$, which encode individual variations in phosphene perception. By training the DSE across a broad distribution of $\phi$ values [23], it learns to produce optimized stimuli tailored to diverse patient profiles.
To further refine these user-specific parameters, we employed human-in-the-loop optimization (HILO), iteratively updating $\phi$ estimates based on pairwise comparisons provided by human participants [23]. Preferences between two candidate parameter sets, $\phi_{1}$ and $\phi_{2}$, were modeled using a Gaussian process:
$P(\phi_{1}\succ\phi_{2}|g)=\Phi\big{(}g(\phi_{1})-g(\phi_{2})\big{)},$ (2)
where $\Phi$ was the normal cumulative distribution, and $g(\phi_{i})$ represented the preference function learned by the Gaussian process. A larger value of $g(\phi_{1})$ relative to $g(\phi_{2})$ indicated a higher likelihood that the participant preferred $\phi_{1}$.
We employed the Maximally Uncertain Challenge acquisition function [24] to balance exploration (identifying uncertain parameter regions) and exploitation (refining the best-known parameters). Specifically:
$\phi_{1}$ $\rightarrow\underset{\phi}{\arg\max}\ \mathbb{E}_{p(g|\mathscr{D})}[g(\phi)],$ (3)
$\phi_{2}$ $\rightarrow\underset{\phi}{\arg\max}\ \mathbb{V}_{p(g|\mathscr{D})}[\Phi(g(\phi)-g(\phi_{1}))],$ (4)
where $\phi_{1}$ was the “champion” (the current best parameter set), and $\phi_{2}$ was the “challenger” (the parameter set for which preferences were most uncertain). Here, $\mathbb{E}$ and $\mathbb{V}$ denoted the expectation and variance, respectively.
Participants guided the optimization process by viewing percepts generated for a target MNIST digit encoded with $\phi_{1}$ and $\phi_{2}$. They selected the percept they perceived as more similar to the target. These pairwise preferences were then used to update the Gaussian process model, iteratively refining the estimates of the participant’s optimal parameters $\phi$.
Figure 3: A) Number of participants who significantly preferred human-in-the-loop optimization (HILO)over the naïve encoder (left) and the deep stimulus encoder (DSE)without HILO (right), based on log odds less than 0 in a linear mixed-effects model. B) Example percepts generated by the HILO encoder, the DSE without HILO, and the naïve encoder for three participants across the three experimental conditions. C) Distribution of log odds for HILO across the three experiments. D) Distribution of the two main user-specific parameters, $\rho$ (phosphene size) and $\lambda$ (axon-aligned elongation), colored by the log odds indicating preference for HILO in the main experiment. E) Median mean squared error (MSE)over the course of optimization for each experiment, with shaded regions denoting the interquartile range (IQR). F) Proportion of duels where participant decisions matched those of the simulated agent.
II-E Experimental Conditions
The study included three experimental conditions, designed to evaluate the performance and robustness of the HILO framework under varying levels of complexity and realism:
  • Main experiment: This was the standard experiment adapted from [23], where the unknown parameters $\phi$ for simulated patients were iteratively optimized using participant feedback.
  • Threshold misspecification (TM): This experiment introduced errors in the assumed threshold amplitudes of up to 300%, testing the framework’s robustness to inaccuracies in user-specific parameters.
  • Out-of-distribution (OOD): In this experiment, the ground-truth $\phi$ values for simulated patients were drawn from a distribution outside the training range of the deep stimulus encoder (DSE), evaluating the framework’s ability to generalize to unseen parameter configurations.
In each condition, the HILO framework was evaluated against two baseline models:
  • Naïve encoder: The approach traditionally used by prostheses, where the target image is reduced to the electrode array resolution, and each pixel’s grayscale value is directly scaled to stimulus amplitude.
  • Deep stimulus encoder (DSE): The same deep stimulus encoder as used in HILO, but instead of user-specific parameters being tuned, they are guessed as the mean of the observed ranges reported in  [23].
III Results
III-A Participant Preferences for HILO vs. Other Encoders
Across all three experimental conditions, participants consistently preferred HILO-optimized stimuli over both the naïve encoder and the non-personalized DSE (Fig. 3A). Example percepts across the three experimental conditions are shown in Fig. 3B. Statistical analysis was performed with a linear mixed-effects logistic regression model, with population and per-subject effects. Across subjects, HILO was significantly preferred over the naïve and DSE encoders for all three experiments ($p<0.0001$). In the main experiment, 16 out of 17 participants favored HILO over the other encoders (defined as having a subject-level log odds of less than 0, indicating preference for HILO). In the threshold misspecification condition, all participants selected HILO. In the out-of-distribution condition, 14 out of 17 participants preferred HILO.
Percepts generated by HILO generally retained greater structure and recognizability compared to baseline methods. In the main experiment, both HILO and DSE produced structured percepts that aligned with target images, whereas the naïve encoder often resulted in highly distorted or unrecognizable shapes. Under threshold misspecification, HILO maintained consistent percepts, whereas DSE outputs varied widely, sometimes producing overexposed phosphenes. In the OOD condition, distortions appeared across all methods, though HILO percepts remained more structured and interpretable than those generated by the alternatives.
The distributions of log odds per participant are shown in Fig. 3C. The mean log odds for HILO over the naïve encoder were $-6.70\pm 0.71$ (main experiment), $-4.57\pm 0.39$ (threshold misspecification), and $-1.95\pm 0.46$ (OOD condition). Lower log odds indicate stronger preference for HILO, confirming its advantage across conditions. Although more participants chose HILO in the threshold misspecification experiment than in the main experiment, their preference was weaker, as reflected in the less negative log odds.
III-B Log Odds Across Individual user-specific Parameters
To assess how individual user-specific factors influenced preference trends, log odds were examined as a function of $\rho$ (phosphene size) and $\lambda$ (phosphene elongation) [9] Fig. 3D. Participants who were assigned larger, more elongated phosphenes exhibited higher log odds in favor of HILO, while those with smaller, minimally distorted percepts showed lower log odds differences between encoders. This pattern suggests that HILO provides the greatest benefit when percepts are highly distorted but offers less improvement when phosphenes are already relatively structured.
III-C Comparison Between Human and Simulated Preferences
Mean squared error (MSE) decreased over the course of optimization (Fig. 3E), confirming that participant choices contributed to improved stimulus encoding. The final loss in the main experiment converged to 0.27, with similar performance in the other two conditions. In contrast, purely simulated experiments from Granley et al.[23], where the goal was to directly minimize MSE, achieved a much lower final loss of 0.07. The slower rate of loss reduction in human trials suggests that participants were not strictly optimizing for pixel-wise reconstruction accuracy, instead incorporating additional perceptual or cognitive factors into their choices.
To further quantify these differences, we compared participant selections to those predicted by the simulated agent from Granley et al.[23] (Fig.3F). Across all three experiments, participants made the same choice as the simulated agent in only about 50% of trials, indicating that human decision-making diverges from the assumptions made in purely simulation-based studies. Despite this variability, participants still overwhelmingly preferred HILO, demonstrating its robustness to real-world human responses and reinforcing the need for empirical validation beyond theoretical models.
IV Discussion
This study demonstrates that human-in-the-loop optimization (HILO)successfully personalizes deep stimulus encoders (DSEs)for individual users, even in the presence of model inaccuracies and noisy human feedback. Sighted participants viewing simulated prosthetic vision consistently preferred HILO-optimized stimuli over both naïve encoding and a standard DSE, reinforcing the viability of this approach for adapting neuroprosthetic devices to individual users. These findings provide critical evidence supporting the application of HILO to real-world neuroprosthetic calibration.
IV-A HILO Stimuli Are Preferred Across Conditions
Across all experimental conditions, participants consistently favored HILO-optimized stimuli over the baseline encoders. Even when the assumed model parameters were misspecified, HILO maintained its advantage, demonstrating robustness to variations in user-specific parameters.
In the out-of-distribution condition, where the DSE was not trained on the tested parameter ranges, a small subset of participants preferred the baseline encoder. This suggests that HILO is most effective when operating within a parameter space it has encountered during training, and its generalization outside of this range may require additional refinements.
Nonetheless, log odds analyses confirmed that HILO still provided a perceptual advantage for most participants, reinforcing its adaptability to individual differences in perception.
IV-B Human Decision-Making Differs from Simulated Patients
Previous studies evaluating HILO [24, 23] relied entirely on simulated patients, whose decisions followed predefined loss functions. While Granley et al. [23] attempted to account for human variability by introducing artificial noise (i.e., demonstrating that HILO remained effective even when simulated patients made random choices in two out of three trials), real human decision-making may not align with these assumptions.
Our results confirm that human choices diverged from simulated predictions. Although MSE decreased over the course of optimization, the final loss was significantly higher than in purely simulated experiments [23], suggesting that human participants did not strictly optimize for pixel-wise accuracy. Additionally, participants selected the same choices as the simulated agent in only about 50% of trials. This suggests that human decision-making incorporates perceptual factors beyond those captured by the computational model.
Despite this divergence, participants still overwhelmingly preferred HILO, reinforcing the need for validation with real users rather than relying solely on theoretical models.
IV-C Limitations and Future Work
While this study provides strong support for the use of HILO in personalizing neuroprosthetic devices, key challenges remain. The next step is testing HILO with blind prosthesis users, but this presents practical and safety considerations. Deep learning models can generate unpredictable outputs, and while neuroprosthetic devices include firmware to enforce safety limits, additional validation is needed to ensure that optimized stimuli do not produce unintended effects. Future work should focus on refining the optimization objective to better align with human perception, possibly incorporating alternative perceptual loss functions that capture high-level visual features rather than pixel-wise accuracy.
Beyond visual prostheses, this approach may generalize to other sensory neuroprosthetic devices, including cochlear implants and tactile feedback systems, where forward models and deep encoders are already in use [27, 28, 29, 30]. By demonstrating HILO’s effectiveness in human participants, this study marks a significant step toward fully individualized neuroprosthetic optimization.